OpenStudy (anonymous):

(x+1)/(x-3)=4-(12/((x^2)-2x-3))

OpenStudy (anonymous):

can you factor the denominaotr of the second fraction: x^2-2x-3

OpenStudy (anonymous):

it is (x-3)(x+1)

OpenStudy (dumbcow):

yes good now add that fraction to other side

OpenStudy (dumbcow):

like this \[\frac{12}{(x-3)(x+1)} + \frac{x+1}{x-3} = 4\]

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

combine fractions multiply 2nd fraction by (x+1)/(x+1) to get common denominator

OpenStudy (anonymous):

and 4 would stay the same?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

k

OpenStudy (anonymous):

would i then remultiply the mononomials?

OpenStudy (dumbcow):

um yes, you mean expand (x+1)^2 now move denominator to right side just as if you are cross-multiplying

OpenStudy (dumbcow):

\[\rightarrow 12 + (x+1)^{2} = 4(x+1)(x-3)\]

OpenStudy (anonymous):

oh my gosh THAANNNKKK YOOOUUUUU

OpenStudy (dumbcow):

your welcome you still need to multiply it out, add like terms and solve the quadratic though :)

OpenStudy (anonymous):

WAIT SO WHAT WOULD THE ANSWER BE ?

OpenStudy (dumbcow):

x = something just multiply it out and get everything to one side

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

The solutions are x = -5/3 x = 5 so you can check your answers

OpenStudy (anonymous):

sweet thanks